# revtex4-1Failed to recognize

### v1

@@ Abstract
Toward experimental implementations.—As our proof is constructive, i.e., we provide a particular Bell inequality that can be violated by a specific (known) quantum implementation, it raises the question whether our results could experimentally be verified for low number of settings $n$. Indeed, the $M_{3322}$ has experimentally already been violated in Christensen et al. (2015), even with qubit systems. Meaning the first interesting case arises for $n=4$. One challenge that an experimental verification has to face is the relatively low violation achievable according to Eq. (38). We argue here that this is a general feature of the family of $M_{nn22}$ inequalities and not particular to the specific implementation used here. To that end, we note that the maximal possible quantum violation of the $M_{nn22}$ inequality decreases generally for all quantum correlations, according to the third level of the NPA hierarchy Navascués et al. (2007). For instance, while it holds in general quantum theory that $M_{3322}(\mathbf{Q})\leq 0.0324$, already for $n=5$ measurements we only can hope to observe a violation of one order of magnitude less, i.e., $M_{5522}(\mathbf{Q})\leq 0.0040$. In contrast to that, by maximizing Eq. (38) we can reach a value of $M_{3322}(\mathbf{Q})\approx 0.0239$ for $n=3$ measurements and $M_{5522}(\mathbf{Q})\approx 0.0035$ for $n=5$ measurements.
@@
Discussion.—We have proven that, for every finite integer $n\geq 2$, there exist quantum correlations that employ genuinely $n$ measurements in any causal theory of nature. That is, there cannot exist a no-signaling theory in which the predictions of quantum theory are reproduced by using only genuinely $n-1$ or less measurements. Astonishingly, given that the predictions for quantum theory on the particular implementation used in our main result are correct, our result implies that nature cannot be described by a finite number of measurements. Our proof relies on an analytic application of the framework of incompatibility structures to the family of $M_{nn22}$, showing that they constitute a device-independent and theory-independent witness of $n$ genuine measurements. We then argued that a practical verification of our results for low $n$, with particular focus on the case $n=4$ are within experimental reach.
Our works proves that any causal description of nature (including quantum and post-quantum theories) does not only require more than $2$ measurements Quintino et al. (2019) but the number of required measurements is actually unbounded. At the same time, our results provide also an analogue to the findings in Kleinmann and Cabello (2016); Kleinmann et al. (2017), which imply that the predictions of quantum theory cannot be reproduced by a finite number of measurement outcomes. There is an interesting parallel between the latter works and our results, as they require the Hilbert space dimension of the quantum system used for the Bell inequality violation to grow with $n$. While it is unclear whether this phenomenon is a necessity (in particular because $M_{3322}$ can be violated with qubits) it is reasonable to assume that it becomes increasingly harder to distinguish quantum theory from a no-signaling theory restricted to $n-1$ measurements, as the relative difference between $n-1$ and $n$ measurements shrinks. High-dimensional quantum systems might therefore be required in general to demonstrate this decreasing gap between the theories.
Except the above question our work also initiates the search for more suitable Bell inequalities and their quantum violation, such that experimental demands in terms of noise-robustness and detector efficiency could be lowered. It would also be interesting to see which other incompatibility structures Kunjwal et al. (2014); Heinosaari et al. (2008); Liang et al. (2011); Quintino et al. (2019) could be ruled out device-independently our even theory-independently. Finally, it might be insightful to consider concepts for incompatibility and entanglement theory simultaneously to rule out possible post-quantum theories.
@@ Acknowledgements.
Supplemental Material for "Nature cannot be described by any causal theory with a finite number of measurements"
In this Supplemental Material, we give a detailed proof to our main result from the main text. In order to make this document self-contained, we quickly review the most necessary concepts and the notation of the manuscript.
We say an assemblage $\mathcal{A}=\{\{A_{a|x}\}_{a}\}_{x}$ is jointly measurable if it can be written such that
$A_{a|x}=\sum_{\lambda}p(a|x\lambda)G_{\lambda},\ \forall\ a,x,$ (39)
holds. Here, $\{G_{\lambda}\}_{\lambda}$ is the so-called parent POVM from which we simultaneously recover the statistics of the whole assemblage by post-processing with the probabilities $p(a|x\lambda)$. For a general assemblage $\mathcal{A}$, we say it contains genuinely $n$ measurements Quintino et al. (2019) if it cannot be decomposed such that
@@
where the set $\mathcal{T}$ contains all tuples $(s,t\neq s)$ constructed from the integers $\{1,\cdots,n\}$. The restriction to the particular incompatibility structure in Eq. (40) leads to a restriction of the achievable behaviours $\mathbf{P}$. More precisely, assume Bob’s measurements obey the above incompatibility structure. For the resulting correlations, it holds that
@@
where $p^{(s,t)}(ab|xy)$ describes a behaviour that is local once it is restricted to only the measurements $(s,t)$ of Bob. Certifying that $\mathbf{P}$ cannot be decomposed as above, proves in a device-independent way that Bob has genuinely access to $n$ measurements. Moreover, if we require $\mathbf{P}$ to only be a no-signaling distribution and not necessarily obey a quantum implementation, it is also a proof of the existence of $n$ genuine measurements on Bob’s side in a theory independent way.
@@
To proof our main result, we will use the Collins-Gisin table representation of Bell inequalities (and behaviours) Collins and Gisin (2004). In the particular case of dichotomic (i.e., two-outcome) measurements, a behaviour $\mathbf{P}$ can uniquely be defined (via normalization and no-signaling constraints) through a table of the form
$\begin{array}[]{c||c c}&P(A_{1})&P(A_{2})\\ \hline\cr\hline\cr P(B_{1})&P(A_{1}B_{1})&P(A_{2}B_{1})\\ P(B_{2})&P(A_{1}B_{2})&P(A_{2}B_{2})\\ \end{array},$ (47)
where $P(A_{1})\coloneqq p_{A}(a=0|x=1)$, $P(A_{1}B_{1})\coloneqq p(a=0,b=0\lvert x=1,y=1)$. The generalization to more than two measurement settings is straightforward. Similarly, we can represent Bell-inequalities via their coefficients in a Collins-Gisin table. For instance, the CHSH inequality Clauser et al. (1969) (in Collins-Gisin notation) given by
@@
$\begin{array}[]{c||c c}&-1&\phantom{-}0\\ \hline\cr\hline\cr-1&\phantom{-}1&\phantom{-}1\\ \phantom{-}0&\phantom{-}1&-1\\ \end{array}\leq 0.$ (52)
With this, we are in the position to state and then proof our main result. Let us denote the set of no-signaling correlations of at most genuinely $n$ measurements according to Eq. (41) by $\mathrm{GNS}_{n}$.
@@
For every $n\geq 2$ there exist quantum correlations $\mathbf{Q}\in\mathrm{GNS}_{n}$ that cannot be reproduced by any no-signaling correlations with at most genuinely $n-1$ measurements, i.e., $\mathbf{Q}\notin\mathrm{GNS}_{n-1}$.
@@
Consider the $M_{nn22}$ Brunner et al. (2006) inequalities given by
@@
Let us first note that the $M_{nn22}$ inequality include the $M_{(n-1)(n-1)22}$ inequality given by
@@
That is, if the inequality $M_{nn22}\leq 0$ is violated, it also follows that the inequality $M_{(n-1)(n-1)22}\leq 0$ has to be violated. This follows from the fact that it holds $P(A_{k})\geq P(A_{k}B_{j})$ and $P(B_{k})\geq P(A_{j}B_{k})$ for any no-signaling distribution. With this, it can be directly seen that the first row of correlations in the Collins-Gisin table $M_{nn22}$ cannot exceed $(n-1)P(B_{1})+P(A_{1})$. Finally, the negative coefficient for the entry corresponding to $p(a=0,b=0|x=n,y=2)$ cannot help to violate the inequality.
In a similar fashion, we can show that Bob’s first measurement in the $M_{nn22}$ inequality has to lead to nonlocal correlations with any other of the remaining $(n-1)$ measurements if only considered pairwise. This can be seen in the following way: We want to show that for measurement $y=1$ and every of the remaining measurements, a version of the CHSH inequality
$\begin{array}[]{c||c c}&-1&\phantom{-}0\\ \hline\cr\hline\cr-1&\phantom{-}1&\phantom{-}1\\ \phantom{-}0&\phantom{-}1&-1\\ \end{array}\leq 0.$ (71)
has to be violated. To that end, note first that the anti-diagonal of $-1$ entries always insures that we can find the corresponding correlations, i.e., the probabilities $p(ab\lvert xy)$ that will appear in the CHSH inequality. Note also, that this always requires including Alice’s first measurement, as this is the only measurement that contributes to the $M_{nn22}$ with a negative marginal. For instance, for measurement $y=1$ and $y=2$, these are given by the table
@@
Now, we only have to make sure that we can also have a marginal $-P(B_{1})$ and similarly $-P(A_{1})$ while still having enough marginal terms with a negative sign left to cancel out positive contributions from the correlation terms. That is, the problem now reduces to a counting problem. It can clearly be seen that the first row has $n$ positive signs, the second row has $n-1$ positive signs and in general, the $k$-th row has $(n-k+1)$ positive signs. Note also, that negative signs for correlations $p(ab\lvert xy)$ can always be ignored, as the only make it harder to violate the inequality. Now, let us use Bob’s marginals to cancel out positive contributions from all columns but the first column. It can readily be verified that using Bob’s marginal, we can cancel out 00 marginals in the $n$-th row, $1$-marginal in the $(n-1)$-st row, $2$ marginals in the $(n-2)$-nd row and so on. The only exception to this pattern is the first row, where we can only cancel out $(n-2)$ instead of $(n-1)$ marginals, as the remaining marginal has to contribute to the CHSH inequality.
Now, this number of available marginals matches exactly the number of positive correlations that need to be cancelled out, where we remember that in the first row, there is a positive correlation that we do not want to cancel out, as it also contributes to the CHSH inequality.
Now we only have to consider the remaining first column. In this column we have $(n-2)$ positive correlations that we need to cancel out. However, this can trivially be done by using the $(n-2)$ negative marginals in Alice’s first measurement that do not contribute to the CHSH inequality. Therefore, in total we see that Bob’s first measurement has to lead to nonlocal correlations (and therefore be incompatible) with any other of his remaining measurements.
In our example above, this means
@@
needs to be violated whenever the $M_{nn22}$ inequality is violated.
Now, as a violation of the $M_{nn22}$ inequality implies a violation of the $M_{(n-1)(n-1)22}$ inequality, we can also directly conclude that Bob’s second measurement (which we regard as hist first in the scenario with only $(n-1)$ measurements) has to lead to nonlocal correlations with any of the remaining measurements. We can repeat this procedure until the last step, where we only consider the CHSH inequality between Bob’s last and second to last measurement.
In conclusion, the violation of the $M_{nn22}$ inequality signifies pairwise incompatibility between any pair of Bob’s measurements. Now, by contra position, if one of the pairs of Bob’s measurements is compatible (and hence leads to local correlations), the $M_{nn22}$ cannot be violated. Hence, the violation of the $M_{nn22}$ inequality signifies genuinely $n$ measurements on Bob’s side. Note that we only used no-signaling arguments to derive that the $M_{nn22}$ inequality signifies genuinely $n$ measurements. That is, we only used that marginals do not have lower probabilities than the correlations they appear in. Therefore, the $M_{nn22}$ inequality signifies genuinely $n$ measurements in any no-signaling theory.
@@
To finish the proof, we need to show that the $M_{nn22}$ inequality can be violated within quantum theory for every $n$. To show that this is indeed the case, we consider the family of states and measurements (defined for any $n$) given in Vértesi et al. (2010), which was developed to find a violation of the $I_{nn22}$ inequality for every $n$. As it turns out, the correlations are strong enough to also work in our case. For completeness, we repeat the construction of Vértesi et al. (2010) here.
@@
It follows that these correlations violate the $M_{nn22}$ inequality up to a value of
$M_{nn22}(\mathbf{Q})=\epsilon^{2}(q_{0}^{2}n-(n-1)),$ (92)

### v2

@@ Abstract
Toward experimental implementations.—As our proof provides specific Bell inequalities that can be violated by a specific (known) quantum implementation, it raises the question whether our results could experimentally be verified for a low number of settings $n$. Indeed, the $M_{3322}$ inequality has experimentally already been violated in Christensen et al. (2015), even with qubit systems. Meaning the first interesting case arises for $n=4$. One challenge that an experimental verification has to face is the relatively low violation achievable according to Eq. (40), resulting in high fidelity requirements. This seems to be a general feature of the family of $M_{nn22}$ inequalities and not particular to the specific implementation used here, as the maximal possible quantum violation of the $M_{nn22}$ inequality decreases generally for all quantum correlations, according to the third level of the NPA hierarchy Navascués et al. (2007). For instance, while it holds in general quantum theory that $M_{3322}(\mathbf{Q})\leq 0.0324$, already for $n=5$ measurements we only can hope to observe a violation of one order of magnitude less, i.e., $M_{5522}(\mathbf{Q})\leq 0.0040$. In contrast to that, by maximizing Eq. (40) we can reach a value of $M_{3322}(\mathbf{Q})\approx 0.0239$ for $n=3$ measurements and $M_{5522}(\mathbf{Q})\approx 0.0035$ for $n=5$ measurements.
Analyzing whether this trend is specific to the family of $M_{nn22}$ inequalities or a generic feature of the considered problem is generally hard, as it requires to find other families of Bell inequalities that are obeyed by all no-signaling distributions of at most genuinely $(n-1)$-wise incompatible measurements, which appears to be a hard task. Moreover, it is known that binary quantum correlations (in arbitrary dimensions) are quite restrictive in their strength, at least when traceless observables are considered Acín et al. (2006).
@@
Discussion.—We have proven that, for every finite integer $n\geq 2$, there exist quantum correlations that employ genuinely $n$-wise incompatible measurements in any no-signaling theory. That is, there cannot exist a no-signaling theory in which the predictions of quantum theory are reproduced by using only convex combinations of genuinely $(n-1)$-wise incompatible measurements across different partitions. If one would claim that such a theory exists, a violation of the corresponding $M_{nn22}$ inequality would falsify said theory. Our proof relies on an analytic application of the framework of incompatibility structures to the family of $M_{nn22}$, showing that it is witness of genuine $n$-wise incompatibility when distributions $\mathbf{P}$ are only restricted by the no-signaling principle. We then argued that a practical verification of our results for low $n$, with particular focus on the case $n=4$ are within experimental reach.
Our works proves that any no-signaling theory that reproduces the predictions of quantum theory does not only require more than $2$ measurements Quintino et al. (2019) but the number of required measurements is actually unbounded. At the same time, our results provide also an analogue to the findings in Kleinmann and Cabello (2016); Kleinmann et al. (2017) (see also Hu et al. (2018) for a corresponding experimental implementation), which imply that the predictions of quantum theory cannot be reproduced by a finite number of measurement outcomes in any no-signaling theory. There is an interesting parallel between the latter works and our results, as they require the Hilbert space dimension of the quantum system used for the Bell inequality violation to grow with $n$. While it is unclear whether this phenomenon is a necessity (in particular because $M_{3322}$ can be violated with qubits) it is reasonable to assume that it becomes increasingly harder to distinguish quantum theory from a no-signaling theory restricted to $(n-1)$ measurements, as the relative difference between $(n-1)$ and $n$ measurements shrinks. High-dimensional quantum systems might therefore be generally required to demonstrate this decreasing gap between the theories.
Except the above question our work also initiates the search for more suitable Bell inequalities and their quantum violation, such that experimental demands in terms of noise-robustness and detector efficiency could be lowered. It would also be interesting to see which other incompatibility structures Kunjwal et al. (2014); Heinosaari et al. (2008); Liang et al. (2011); Quintino et al. (2019); Yadavalli et al. (2024) could be ruled out on the level of the no-signaling assumption only. Finally, it might be insightful to consider concepts for incompatibility and entanglement theory simultaneously to rule out possible post-quantum theories.
@@ Acknowledgements.
Appendix for "Quantum correlations cannot be reproduced with a finite number of measurements in any no-signaling theory"
In this Appendix, we give a detailed proof of our main result from the main text. In order to make this document self-contained, we repeat the necessary concepts and the notation of the manuscript and provide aiding examples.
We consider quantum measurements, which are described by positive semidefinite operators $A_{a|x}\geq 0$ acting on a Hilbert space $\mathcal{H}\cong\mathds{C}^{d}$ of dimension $d$, such that the completeness relation $\sum_{a}A_{a|x}=\mathds{1}$, (where $\mathds{1}$ is the identity operator), is fulfilled. Here, $a$ denotes the outcome corresponding to the measurement $x$. We say an assemblage $\mathcal{A}=\{\{A_{a|x}\}_{a}\}_{x}$ is jointly measurable if it can be written such that
$A_{a|x}=\sum_{\lambda}p(a|x,\lambda)G_{\lambda},\ \forall\ a,x,$ (41)
holds. Here, $\{G_{\lambda}\}_{\lambda}$ is the so-called parent POVM from which we simultaneously recover the statistics of the whole assemblage by post-processing with the probabilities $p(a|x\lambda)$. Eq. (41) can be understood as checking whether the assemblage $\mathcal{A}$ can be seen as one effective measurement (via the parent POVM). However, if Eq. (41) cannot be satisfied for a given assemblage $\mathcal{A}$, it does not guarantee that $\mathcal{A}$ cannot be reduced to effectively two or more measurements. To do so, we have to make use of more general incompatibility structures and the notion of genuine $n$-wise incompatible assemblages.
For a general assemblage $\mathcal{A}$, we say it is genuine $n$-wise incompatible (contains genuinely $n$ measurements) Quintino et al. (2019) if it cannot be decomposed such that
@@
where $\mathcal{F}^{(s,t)}\in\mathrm{JM}^{(s,t)}$ are assemblages, employed with probability $p_{(s,t)}$, in which the pair $(s,t\neq s)$ is jointly measurable. This makes each $\mathcal{F}^{(s,t)}$ contain at most genuinely $(n-1)$-wise incompatible measurements (as the pair $(s,t\neq s)$ is one effective measurement) and the set $\mathcal{T}$ contains all tuples $(s,t\neq s)$ constructed from the integers $\{1,\cdots,n\}$. As an illustrative example, let us consider the incompatibility structures with respect to the three Pauli observables given by the Pauli matrices $X,Y,Z$, also used in Quintino et al. (2019). More precisely, let $\Pi_{a|x}$ with $a\in\{\pm 1\}$ and $x\in\{X,Y,Z\}$ be the projectors corresponding to the Pauli observables. Moreover, consider a noise model of the form
$M^{\eta}_{a|x}=\eta\Pi_{a|x}+(1-\eta)\dfrac{\mathds{1}}{2},$ (43)
with $0\leq\eta\leq 1$. It follows that all Pauli measurements are jointly measurable if $\eta\leq\tfrac{1}{\sqrt{3}}$ and each pair of measurements is compatible for $\eta\leq\tfrac{1}{\sqrt{2}}$. However, the Pauli measurements are only genuinely triplewise incompatible (can be regarded genuinely as $3$ measurements) if $\eta>\tfrac{\sqrt{2}+1}{3}$.
The restriction to the particular incompatibility structure in Eq. (42) leads to a restriction of the achievable behaviors $\mathbf{P}$. More precisely, let us assume a bipartite Bell scenario in which Bob’s measurements obey the above incompatibility structure. It holds for the resulting correlations that
@@
where $\{p^{(s,t)}(ab|xy)\}_{a,b,x,y}$ describes a behavior $\mathbf{P}^{(s,t)}$ that is local once it is restricted to only the measurements $(s,t)$ on Bob’s side. Certifying that $\mathbf{P}$ cannot be decomposed as above, proves in a device-independent way that Bob has genuinely access to $n$ measurements. Moreover, if we require $\mathbf{P}$ to only be a no-signaling distribution and not necessarily obey a quantum implementation, it is also a proof of the existence of genuine $n$-wise incompatible measurements on Bob’s side in any no-signaling theory.
@@
To prove our main result, we will use the Collins-Gisin table representation of Bell inequalities (and behaviors) Collins and Gisin (2004). In the particular case of dichotomic (i.e., two-outcome) measurements, a behavior $\mathbf{P}$ can uniquely be defined (via normalization and no-signaling constraints) through a table of the form
$\begin{array}[]{c||c c c}&P(A_{1})&P(A_{2})&P(A_{3})\\ \hline\cr\hline\cr P(B_{1})&P(A_{1}B_{1})&P(A_{2}B_{1})&P(A_{3}B_{1})\\ P(B_{2})&P(A_{1}B_{2})&P(A_{2}B_{2})&P(A_{3}B_{2})\\ P(B_{3})&P(A_{1}B_{3})&P(A_{2}B_{3})&P(A_{3}B_{3})\end{array},$ (51)
where $P(A_{1})\coloneqq p_{A}(a=0|x=1)$, $P(A_{1}B_{1})\coloneqq p(a=0,b=0\lvert x=1,y=1)$. The generalization to more than three measurement settings is straightforward. Similarly, we can represent Bell-inequalities via their coefficients in a Collins-Gisin table. For instance, the CHSH inequality Clauser et al. (1969) (in Collins-Gisin notation) given by
@@
$\mathrm{CHSH}=$ $\begin{array}[]{c||c c}&-1&\phantom{-}0\\ \hline\cr\hline\cr-1&\phantom{-}1&\phantom{-}1\\ \phantom{-}0&\phantom{-}1&-1\\ \end{array}\leq 0.$ (56)
To compute the corresponding CHSH value of a behavior, we then write $\mathrm{CHSH}\odot\mathbf{P}$, which corresponds to evaluating the CHSH functional for the distribution $\mathbf{P}$. With this, we are in the position to state and then proof our main result. Let us denote the set of no-signaling correlations of at most genuinely $(n-1)$-wise incompatible measurements by $\mathrm{GNS}_{n-1}$. That is, $\mathbf{P}\in\mathrm{GNS}_{n-1}$ if a decomposition according to Eq. (44) exists.
@@
For every $n\geq 2$ there exist quantum correlations $\mathbf{Q}$ obtained from performing $n$ quantum measurements that cannot be reproduced by any no-signaling correlations with at most genuinely $(n-1)$-wise incompatible measurements, i.e., $\mathbf{Q}\notin\mathrm{GNS}_{n-1}$.
@@
Consider the $M_{nn22}$ Brunner et al. (2006) inequality given by
@@
Let us first note that the $M_{nn22}$ inequality includes the $M_{(n-1)(n-1)22}$ inequality given by
@@
That is, if $M_{nn22}\odot\mathbf{P}>0$ for some no-signaling behavior $\mathbf{P}$, it also follows that $M_{(n-1)(n-1)22}\odot\mathbf{P}>0$. This follows from the fact that it holds $P(A_{k})\geq P(A_{k}B_{j})$ and $P(B_{j})\geq P(A_{k}B_{j})$ for any no-signaling distribution, as marginal probabilities have to be at least as large as the correlation terms they appear in. With this, it can be directly seen that the first row of correlations in the Collins-Gisin table $M_{nn22}$ cannot exceed $(n-1)P(B_{1})+P(A_{1})$. Finally, the negative contribution of the entry corresponding to $p(a=0,b=0|x=n,y=2)$ cannot help to violate the inequality. More formally, we use that it holds
$\large[\sum_{j=2}^{n}P(A_{j}B_{1})-(n-1)P(B_{1})\large]+[P(A_{1}B_{1})-P(A_{1})]-P(A_{n}B_{2})\leq 0,$ (72)
for any no-signaling behavior $\mathbf{P}$.
We will use this property as follows. We will first show that in order for the $M_{nn22}$ inequality to be violated, it requires that Bob’s first setting, i.e., $y=1$ violates the CHSH inequality together with any other of his measurements $y\neq 1$. Now, the same has to be true for Bob’s setting $y=2$, as it becomes the first setting of the $M_{(n-1)(n-1)22}$ inequality. Then, by induction, any of Bob’s measurements $y=s$ has to be incompatible with any other setting $t\neq s$. Finally, by the same contra position argument as in the main text and using the linearity of the $M_{nn22}$ inequality together with fact that the decomposition in Eq.(44) is convex, we know that it holds $M_{nn22}\odot\mathbf{R}\leq 0$ for any behavior $\mathbf{R}\in\mathrm{GNS}_{n-1}$ according to Eq. (46).
To show that Bob’s first measurement in the $M_{nn22}$ inequality has to lead to nonlocal correlations (and hence be incompatible) with any other of the remaining $(n-1)$ measurements if only considered pairwise we proceed in the following way. We want to show that for measurement $y=1$ and every of the remaining measurements $y\neq 1$, a version of the CHSH inequality
$\mathrm{CHSH}=$ $\begin{array}[]{c||c c}&-1&\phantom{-}0\\ \hline\cr\hline\cr-1&\phantom{-}1&\phantom{-}1\\ \phantom{-}0&\phantom{-}1&-1\\ \end{array}\leq 0.$ (76)
has to be violated. This can be seen from counting arguments using the specific Collins-Gisin table representation of the $M_{nn22}$ inequality in Eq. (64). Note first that the anti-diagonal of $-1$ entries in the formulation of $M_{nn22}$ always ensures that we can find the corresponding coefficients for the correlation terms, i.e., the probabilities $p(ab\lvert xy)$ will appear with the same sign in the CHSH inequality. Note further, that this always requires including Alice’s first measurement, as this Alice’s only measurement that contributes to the $M_{nn22}$ with a negative marginal. For instance, for measurement $y=1$ and $y=2$, these correlation terms are given by the table
@@
Of course, we also have to include Alice’s and Bob’s marginal corresponding to the CHSH inequality. That is, we need to show that the inequality
@@
can be extracted from the $M_{nn22}$ inequality by showing that the remaining terms will contribute non-positively towards the Bell value for any no-signaling distribution $\mathbf{P}$.
Let us first list the following facts about the Collins-Gisin table corresponding to the $M_{nn22}$ inequality:
  1. 1.
    In row $r\in[1,n]$ there are $(n+1-r)$ positive ($+1$) valued correlation terms.
  2. 2.
    In row $r$, there are $(n-r)$ negative ($-1$) marginals of Bob.
  3. 3.
    In column $1$, there are $n$ positive ($+1$) correlation terms.
  4. 4.
    In column $1$, there are $(n-1)$ negative ($-1$) marginals of Alice.
With that, we proceed with the following algorithmic construction to show that $M_{nn22}\odot\mathbf{P}>0$ implies a violation of the CHSH inequality using Bob’s measurements $y_{1}=1$ and $y_{2}=s$ for every $s\in[2,n]$. For a given $s$, let us fix Alice’s measurements $x_{1}=1$ and $x_{2}=n-s+2$. The following algorithm will extract the CHSH inequality for the settings $x_{1},x_{2},y_{1},y_{2}$ while only removing non-positive terms for any no-signaling distribution $\mathbf{P}$.
  1. 1.
    In row $y_{1}=1$, cancel out all $+1$ correlation terms except the term corresponding to $x_{1}$ and $x_{2}$, using $(n-2)$ of the negative marginals of Bob. This fixes the correct entries in the first row. That means, in the first row, the only remaining terms correspond to $-P(B_{y_{1}})$, $P(A_{x_{1}}B_{y_{1}}),$ and $P(A_{x_{2}}B_{y_{1}})$.
  2. 2.
    In row $y_{2}=s$, cancel out all $+1$ correlation terms except the term corresponding to $x_{1}=1$ using the $(n-s)$ negative marginal terms of Bob. This leaves only one remaining term, corresponding to $P(A_{x_{1}}B_{y_{2}})$.
  3. 3.
    Set all $-1$ correlation terms except the term corresponding to $P(A_{x_{2}}B_{y_{2}})$ to 00.
  4. 4.
    In all rows $r$, except $y_{1},y_{2}$, cancel out all $+1$ correlation terms except those corresponding to column $x_{1}=1$ using Bob’s $(n-r)$ negative marginals.
  5. 5.
    Cancel out all $+1$ correlation terms in column $x_{1}=1$, except those in row $y_{1},y_{2}$, using $(n-2)$ of the negative marginal terms of Alice, leaving one term corresponding to $-P(A_{x_{1}})$ remaining.
It follows directly, that the above algorithm picks out the CHSH inequality between the settings $x_{1},x_{2},y_{1},y_{2}$, while only removing non-positive terms in each step, as marginals outweigh their corresponding correlation terms and probabilities are positive. Hence, we have shown the required property that none of Bob’s measurements can be pairwise compatible when the $M_{nn22}$ inequality is violated. By the above reasoning, this directly implies $M_{nn22}\odot\mathbf{R}\leq 0$ for any no-signaling distribution $\mathbf{R}\in\mathrm{GNS}_{n-1}$.
@@
To finish the proof, we need to show that the $M_{nn22}$ inequality can be violated within quantum theory for every $n$. To show that this is indeed the case, we consider the family of states and measurements (defined for any $n$) given in Vértesi et al. (2010), which was developed to find a violation of the $I_{nn22}$ inequality for every $n$. As it turns out, the correlations are strong enough to also work in our case. For completeness, we repeat the construction of Vértesi et al. (2010) here in the following.
Before giving the explicit quantum strategy, we would like to point out the construction found in Vértesi et al. (2010) is tailored to a different (but equivalent) representation of the $I_{nn22}$ (and therefore the resulting $M_{nn22}$) inequality then used in the remainder of this work. In particular, the quantum strategy down below can be used to violate the $\tilde{M}_{nn22}$ given by
$\tilde{M}_{nn22}=$ $-(n-1)P(A_{1})-\sum_{y=2}^{n}P(B_{y})+\sum_{y=1}^{n}P(A_{1}B_{y})$ (93)
$+\sum_{x=2}^{n}P(A_{x}B_{x})-\sum_{1\leq y<x\leq n}P(A_{x}B_{y})\leq 0,$
which in Collins-Gisin table notation is given by
$\tilde{M}_{nn22}=\begin{array}[]{c|| ccccccc}\mbox{}&-(n-1)&\phantom{-}0&\phantom{-}0&\phantom{-}0&\cdots&\phantom{-}0&\phantom{-}0\\ \hline\cr\hline\cr\phantom{-}0&\phantom{-}1&-1&-1&-1&\cdots&-1&-1\\ -1&\phantom{-}1&\phantom{-}1&-1&-1&\cdots&-1&-1\\ -1&\phantom{-}1&\phantom{-}0&\phantom{-}1&-1&\cdots&-1&-1\\ -1&\phantom{-}1&\phantom{-}0&\phantom{-}0&\phantom{-}1&\cdots&-1&-1\\ \phantom{-}\vdots&\phantom{-}\vdots&\phantom{-}\vdots&\phantom{-}\vdots&\mbox{}&\ddots&\mbox{}&\phantom{-}\vdots\\ -1&\phantom{-}1&\phantom{-}0&\phantom{-}0&\phantom{-}0&\cdots&\phantom{-}1&-1\\ -1&\phantom{-}1&\phantom{-}0&\phantom{-}0&\phantom{-}0&\cdots&\phantom{-}0&\phantom{-}1\end{array}\leq 0.$ (102)
However, the inequalities ${M}_{nn22}\leq 0$ and $\tilde{M}_{nn22}\leq 0$ are simply related by symmetries that do not alter local bounds or any achievable quantum violations. In particular, after exchange of Alice’s settings (corresponding to exchanging columns) $x\leftrightarrow n-x+2$ for all $x\geq 2$ and exchanging Alice’s outcomes $a=0\leftrightarrow a=1$ for all $x\geq 2$, which amounts to replacing $P(A_{x}B_{y})$ with $P(B_{y})-P(A_{x}B_{y})$ and $P(A_{x})$ with $1-P(A_{x})$, we obtain the form of the ${M}_{nn22}$ inequality used in Eq. (64). Hence, the following proof for a quantum violation of $\tilde{M}_{nn22}\leq 0$ translates directly to a quantum violation of $M_{nn22}\leq 0$ with the same Bell score.
@@
It follows that these correlations violate the $\tilde{M}_{nn22}$ inequality up to a value of
$\tilde{M}_{nn22}(\mathbf{Q})=\epsilon^{2}(q_{0}^{2}n-(n-1)),$ (107)